Planning week 23-30 January 2010

The week began with AR 11041 being the only active region on the Sun, in the SE quadrant. On 20 January it had produced two M-flares while it was at the limb, however, by the start of the planning period (11 UT, 23 Jan) the region was very quiet (few B-flares). Another active region was present in the NW quadrant for the week, but no observations were performed on this.

The week was divided into three OP periods: OP1 was 3 days long, and OP2 and OP3 were 2 days long.

Key observations

  1. From Saturday to Tuesday I followed AR 11041 from 10:45 on 23 Jan to 18:00 on 25 Jan, with a 5 hour gap on 24 Jan from 18:00 to 23:00, when HOP 146 was supported.
  2. HOP 146 was supported twice on 24 Jan 18:00-23:00 and 25 Jan 18:00-23:00
  3. J. Klimchuk nanoflare study (sit-and-stare) was performed on 26 and 27 January at two pointings within AR 11041.
  4. HOP 100 was supported with a large diagnostic map on 26-27 January.
  5. A slit calibration study was performed at the east limb.
  6. Three different types of AR outflow study were performed for 28-30 Jan on the east footpoint regions of AR 11041.
  7. HOP 79 (N-S irradiance scan) was supported on 28-29 Jan.
  8. HOP 81 (South coronal hole) was supported on 29 Jan.

Saturday 23 January (11-24 UT)

EIT image at 13:13 UT
GOES plot (23-26 Jan)

AR 11041 monitored continuously with FLR001_flare_study, with PRY_slot_context_v3 interspersed every 6 hours or so. I slipped in a FILL001 study at 17:43 to get a spectral atlas:

10:47    PRY_slot_context_v3
10:50    FLR001_flare_study (x85)
17:39    PRY_slot_context_v3
17:43    FILL001
18:06    PRY_slot_context_v3
18:09    FLR001_flare_study (x73)

SOT plan (OP1)
XRT plan (OP1)

Sunday 24 January 

EIT image at 11:48 UT

AR 11041 monitoring with FLR001_flare_study continued until 06:00:

00:00    PRY_slot_context_v3
00:04    FLR001_flare_study (x72)
05:51    PRY_slot_context_v3

For the SAA-free period I scheduled a slot movie (PRY_slot_context_v3) from 10:30 for 4hrs 40mins. The movie from this study shows some activity in the core of the AR.

At 18:00 HOP 146 (coronal holes; PI: Ko) was supported at the south pole where there was quite a nice coronal hole.

Monday 25 January

EIT image at 18:24 UT

At 06:46 I slipped in a Eclipse_raster_v2 study to get good spectral coverage of AR 11041. For the SAA-free period I ran cam_artb_lite_v2 for 4 hours, from 09:37.

At 18:00 HOP 146 (coronal holes; PI: Ko) was supported at the north pole where the coronal hole was very small.

Tuesday 26 January (00-11 UT)

No observations due to telemetry restrictions.

Tuesday 26 January (11-24 UT)

EIT image at 11:48 UT
GOES plot, 26-28 Jan

A nanoflare study for J. Klimchuk was run from 11:01 to 12:15:

11:01    PRY_slot_context_v3
12:14    PRY_slot_context_v3

The pointing was set to the east side of the AR at the footpoints of hot loops (moss). A 2nd pointing was performed on 27 Jan (see below). Following this study, single runs of PRY_slot_context_v3 were scheduled every hour before the start of HOP 100.

HOP 100 was supported from 18:12 onwards. SOT takes a large SP map for this, requiring multiple satellite pointings. I requested XRT to run a DEM study, but apparently they're restricted to three filters at the moment. I decided to run a large diagnostic map with EIS. Note that the SOT data were taken between 21:20 and 05:00. The timeline for the observations is

18:12    PRY_slot_contextLITE (x 7)
18:57    HPW017AR_30x400_30s2 (x4, stitching)
20:34    HPW017AR_30x400_30s2 (x4, stitching)
22:11    HPW017AR_30x400_30s2 (x4, stitching)
23:42    HPW017AR_30x400_30s2 (x2, stitching)
00:35    HPW017AR_30x400_30s2 (x1)
01:09    HPW017AR_30x400_30s2 (x2, stitching)

Before each of the HPW017 studies there is a PRY_slot_contextLITE context raster.

The GOES plot shows a number of B-flares between 18:00 and 06:00, although these may be from the active region in the NW quadrant.

SOT plan (OP2)
XRT plan (OP2)

Wednesday 27 January

EIT image at 16:25 UT

HOP 100 support was continued with some slot context movies:

01:52    PRY_slot_contextLITE (x 10)
02:42    PRY_slot_contextLITE (x 22)
04:20    PRY_slot_contextLITE (x 13)
05:02    PRY_slot_contextLITE (x 9)

The second pointing (west side of AR) of the Klimchuk nanoflare study was performed between 07:35-08:50:

07:35    PRY_slot_context_v3
08:48    PRY_slot_context_v3

Three further PRY_slot_context_v3 studies were run at 09:12, 10:15 and 11:15.

No further studies were run except the SYNOP.

Thursday 28 January (00-11 UT)

A slit alignment study was run at the east limb, which is very quiet:

05:53    SK_DEEP_5x512_SLIT1 (top)
06:04    SK_DEEP_5x512_SLIT1 (bottom)
06:15    SK_DEEP_10x512_SLIT2 (top)
06:20    SK_DEEP_10x512_SLIT2 (bottom)

Thursday 28 January (11-24 UT)

TRACE 171 mosaic at 00:16 UT
GOES plot 28-30 Jan

For the period 10:37-15:23 I'm doing an AR outflow study. I first take a detailed raster (PRY_footpoints_HI2, 10:37) at the east footpoints and then a slot context movie (PRY_slot_context_v3, 11:19). The idea is to prove if any flows seen in the context movie actually correspond to blueshifts.

Between 15:30 and 06:00 EIS supports HOP 79 doing a N-S scan.

SOT plan (OP3)
XRT plan (OP3)

Friday 29 January

TRACE 171 mosaic at 00:16 UT

HOP 79 is continued to 06:00 today (see above).

For the period 07:10-08:22 I continue the AR outflow study. This time I do the detailed raster (PRY_footpoints_HI2, 07:13) followed by four runs of SK_DEEP_10x512_SLIT2. I position these in pairs such that the first pair covers the top and bottom of the CCD in the proposed outflow region, just to the left of the footpoints. The second pair is positioned 20 arcsec to the left, so further out in the outflow region.

The pointing for PRY_footpoints_HI2 looks like it might be good: see the over-plot on a TRACE 171 image.

HOP 81 (south polar coronal hole) is supported between 10:26 and 15:23.

Between 16:00 and 06:05 I do further AR outflow studies. Firstly, for 16:00-19:09 I do a slot context movie (PRY_slot_context_v3). I then do a high cadence, small raster study with GDZ_LT_PLUME_10x180 for the period 19:25-06:00.

The pointing for the GDZ study was not ideal: see this over-plot on a TRACE 171 image.

Saturday 30 January (0-11 UT)

TRACE 171 mosaic at 00:26 UT

The AR outflow study from 29 January is continued until 06:00 (see above).